OpenStudy (anonymous):

suppose that y varies inversely with x. write a function that models the inverse function. x=15 when y=2

OpenStudy (vispla22):

the equation is y=k/x

OpenStudy (vispla22):

plug the values in and solve for k 2=k/15

OpenStudy (vispla22):

k=30

OpenStudy (vispla22):

the equation is y=30/x

OpenStudy (anonymous):

thanks guys!